# Dependency Parsing is More Parameter-Efficient with Normalization

### v1

@@ 4.1 Data
To train and evaluate our models, we use four SemDP datasets, [^1] We neglect ACE04 and ACE05, two popular datasets for relation extraction, due to their prohibitive cost. along with the 2.2 version of the Universal Dependencies English EWT treebank (enEWT) ud22 and SciDTB yang2018scidtb .
@@
As regards SemDP, ADE gurulingappa2012ade is a medical-domain dataset comprising reports of drug adverse-effect reactions. Each disease is associated to a drug which caused it through the only type of relation present, i.e. “adverseEffect.” CoNLL04 roth2004conll04 contains news texts and is annotated with the classic entities $e_{i}\in$per, org, loc} and relations $r_{j}\in$workFor, kill, orgBasedIn, liveIn, locIn}. ADE and CoNLL04 are characterized by relations which are not complex enough to form connected graphs of considerable size. SciERC luan2018scierc is a dataset compiled from sentences extracted from literature in the domain of artificial intelligence. It is arguably the most challenging dataset used in this study, since most of the entities in the validation and testing partitions are not assigned an entity class. This makes it hard to train tag embeddings that can help to infer edges. ERFGC yamakata_english_2020 is a dataset comprising “flow graphs” parsed from culinary recipes. The semantic dependency graphs of these recipes are directed and acyclic, with a single root. Note that, differently from bhatt_end–end_2024 , and as advised directly by yamakata_english_2020 in personal correspondence, we ignore the “-” edge labels present in the corpus.
@@
As regards SynDP, we use enEWT ud22 to compare directly against ji_graph-based_2019 ’s results. While many works evaluate SynDP on the Penn Treebank prasad-etal-2008-penn it is closed access and prohibitively expensive. Instead, we use SciDTB yang2018scidtb , a discourse analysis dataset comprising 798 abstracts extracted from the ACL Anthology. It was processed for the syntax dependency parsing task using Stanza qi2020stanza . From these two datasets, we consider the xPOS tags (e.g. noun, preposition, determiner) and the syntactic dependencies (e.g. sentence root, object, adverbial modifier) to respectively be the entities and relations of the dependency graphs to infer.
For ADE, CoNLL04, and SciERC we use the splits provided in bi_codekgc_2024 . [^2] https://drive.google.com/drive/folders/1vVKJIUzK4hIipfdEGmS0CCoFmUmZwOQV ERFGC is not available online; we obtained it by contacting the authors of yamakata_english_2020 . Table 1 summarizes the statistics of the datasets used in this work and reports their entity and relation class annotations.
@@ 4.2 Model
$\textbf{h}_{i}^{tag}=~{}$ $\phi(\textbf{x}_{i}),\quad\textbf{h}^{tag}_{i}\in\mathbb{R}^{d_{h}}$
$\textbf{y}_{i}^{tag}=~{}$ $\text{Softmax}(\text{MLP}^{tag}(\textbf{h}_{i}^{tag})),\quad\textbf{y}_{i}^{tag}\in\mathbb{R}^{\mid T\mid}$
@@
In the parser, an optional $k$-layered BiLSTM $\psi$ produces new representations $\textbf{h}_{i}=\psi(\textbf{e}_{i}^{tag}\oplus\textbf{x}_{i})$, which are then projected into four different representations:
$\textbf{e}_{i}^{h}=~{}\text{MLP}^{(edge-head)}(\textbf{h}_{i}),~{}\textbf{e}_{i}^{d}=~{}\text{MLP}^{(edge-dept)}(\textbf{h}_{i})$
$\textbf{r}_{i}^{h}=~{}\text{MLP}^{(rel-dept)}(\textbf{h}_{i}),~{}~{}~{}~{}\textbf{r}_{i}^{d}=~{}\text{MLP}^{(rel-head)}(\textbf{h}_{i})$
@@
$f(\textbf{x}_{1},\textbf{x}_{2};W)=\textbf{x}_{1}^{\top}W\textbf{x}_{2}+\textbf{x}_{1}^{\top}\textbf{b}$
$s_{i}^{edge}$ $=f^{(edge)}(\textbf{e}_{i}^{h},\textbf{e}_{i}^{d};W_{e}),\quad W_{e}\in\mathbb{R}^{d\times 1\times d}$
$s_{i}^{rel}$ $=f^{(rel)}(\textbf{r}_{i}^{h},\textbf{r}_{i}^{d};W_{r}),\quad W_{r}\in\mathbb{R}^{d\times\mid R\mid\times d}$
where $R$ is the set of relation classes, i.e. the possible labels applied to an edge.
Finally, in the decoder, the edge scores are used in conjunction with the relation representations $\textbf{r}_{i}^{h}$ and $\textbf{r}_{i}^{d}$ to obtain the final predictions. During training, we do greedy decoding, while during inference, we use Chu-Liu/Edmonds’ maximum spanning tree (MST) algorithm edmonds_optimum_1967 to ensure the predictions are well-formed trees. This is especially useful with big dependency graphs, since greedy decoding is more likely to produce invalid trees as size increases. When doing greedy decoding, an edge index (i.e. an adjacency matrix) $a_{i}=\arg\max_{j}s_{ij}^{edge}$ is produced by taking the argmax of the attention scores $s_{i}^{edge}$ across the last dimension. The edge index is then used to select which head relation representations $r_{i}^{h}$ to use to calculate the relation scores $s_{i}^{rel}=f(\textbf{r}_{i}^{h},\textbf{r}_{i}^{d};W),\quad W\in\mathbb{R}^{d\times\mid R\mid\times d}$. The relations are then predicted as $r_{i}=\arg\max_{j}s_{ij}^{rel}$. When using MST decoding, edge and relation scores are combined into a single energy matrix where each entry represents the score of a specific head-dependent pair with its most likely relation type. This energy matrix is then used in the MST algorithm, producing trees with a single root and no cycles. Prior to energy calculation, edge scores and relation scores are scaled so that low values are squished and high values are increased, making the log softmax produce a hard adjacency matrix.
@@
$\mathcal{L}_{tag}$ $=-\frac{1}{|\mathcal{V}|}\sum_{i=1}^{|\mathcal{V}|}\sum_{t=1}^{|T|}y_{i,t}^{tag}\,\log p\bigl{(}y_{i,t}^{tag}\bigr{)}$
$\mathcal{L}_{edge}$ $=-\sum_{i,j=1}^{|\mathcal{V}|}\log p\bigl{(}y_{i,j}^{edge}=1\bigr{)}$
$\mathcal{L}_{rel}$ $=-\sum_{i=1}^{|\mathcal{V}|}\sum_{j=1}^{|\mathcal{V}|}\mathbbm{1}\bigl{(}y_{i,j}^{edge}=1\bigr{)}\sum_{\ell=1}^{|R|}y_{i,j,\ell}^{rel}\,\log p\bigl{(}y_{i,j,\ell}^{rel}\bigr{)}$
$\mathcal{L}$ $=\lambda_{1}\,\mathcal{L}_{tag}+\lambda_{2}\bigl{(}\mathcal{L}_{edge}+\mathcal{L}_{rel}\bigr{)}\,$
Losses are calculated based on the gold tags, edges, and relations. We set $\lambda_{1}=0.1$ and $\lambda_{2}=1$ as hyperparameters because the tagging task is much simpler than predicting the edges, since the same top performance is always achieved regardless of any other selected architecture hyperparameters. Following the usual approach for SynDP dozat_deep_2017 ; ji_graph-based_2019 ; jiang_entity-relation_2024 , when training on enEWT and SciDTB we use an oracle, the gold tags, and do not predict the POS tags ourselves. Since in this case we only focus on training the edge and relation classification tasks, we set $\lambda_{1}=0$.
@@ 4.3 Hyperparameters
We experiment with a range of hyperparameters for the encoder, tagger, and parser, as listed in Table 2. We use BERTbase devlin-etal-2019-bert as our pre-trained encoder, which we keep frozen throughout the whole training run in our main setting. As regards the tagger, we set $L_{\phi}=1$ and $h_{\phi}=100$, as in bhatt_end–end_2024 , with weights initialized with a Xavier uniform distribution. In Appendix B.1, we ablate the $\phi$ BiLSTM and the use of the tag embeddings $\mathbf{e}^{tag}_{i}$ to assess their impact on overall performance. Finally, with relation to the parser, for all hyperparameter combinations of $\{L_{\psi},h_{\psi},d_{\text{MLP}}\}$, we run our experiments by initializing its weights with a Xavier uniform distribution ($I_{par}=\mathcal{U}$) and no LayerNorm $\texttt{LN}_{\psi}$. In Sections B.2, B.3, B.4 and B.5, we conduct ablations over the parser hyperparameters indicated in Table 2.
Table 2: Hyperparameter ranges tested in our experiments. $\psi$ = Parser BiLSTM. $f$ = biaffine layer.
@@
MLP(edge) and $f^{(edge)}$ init. $I_{\text{par}}\in\{\mathcal{U},\mathcal{N}\}$
@@
We train our models for $2k$ steps and evaluate on the development partition of each dataset every 100 steps for ADE, CoNLL04, SciERC, and ERFGC. For enEWT and SciDTB, we train for $5k$ steps with $1k$-step validation intervals to make our results more comparable with the start of the art ji_graph-based_2019 . We apply early stopping at 30% of the total steps without improvement and choose the best model based on top performance on the development split. In Appendix B.6, we extend the training to $10k$ steps and fully fine-tune a variety of small and large pre-trained language models to show time-wise test performance trends more clearly. We set the learning rate at $\eta=1\times 10^{-3}$ when the encoder is kept frozen. In all settings, including ablations, we use AdamW loshchilov2019adamw as the optimizer and a batch size of 8.
@@ 4.4 Evaluation
Following bhatt_end–end_2024 ; dozat_simpler_2018 ; jiang_entity-relation_2024 , we measure tagging and parsing performance on ADE, CoNLL04, SciERC, and ERFGC in terms of micro-averaged F1-measure. In addition, for enEWT and SciDTB we use unlabeled (UAS) and labeled (LAS) attachment score nivre2017universal . To corroborate the validity of our results, we train and evaluate each setting, including ablations, with five random seeds. We report mean and standard deviation for the F1, UAS, and LAS metrics, averaged over the five runs. For brevity, the performance on the tagging and unlabeled edge prediction tasks are reported in Appendix A. To test the significance of our results, we use the one-tailed Wilcoxon signed-rank test woolson2005wilcoxon .
@@ 5 Results and discussion
Model $a$ $L_{\psi}$ ADE CoNLL04 SciERC ERFGC enEWT SciDTB
bhatt_end–end_2024 $1$ 3 0.653±$\pm 0.018$ 0.566±$\pm 0.019$ 0.257±$\pm 0.024$ 0.701±$\pm 0.009$ 0.804 ±$\pm 0.006$ 0.915 ±$\pm 0.002$
Ours $1$ 0 0.541 ±$\pm 0.021$ 0.399 ±$\pm 0.024$ 0.147 ±$\pm 0.049$ 0.548 ±$\pm 0.010$ 0.559 ±$\pm 0.005$ 0.729 ±$\pm 0.004$
1 0.657 ±$\pm 0.011$ 0.556 ±$\pm 0.021$ 0.282 ±$\pm 0.009$ 0.676 ±$\pm 0.010$ 0.771 ±$\pm 0.006$ 0.892 ±$\pm 0.002$
2 0.667 ±$\pm 0.011$ 0.573 ±$\pm 0.025$ 0.273 ±$\pm 0.010$ 0.694 ±$\pm 0.010$ 0.796 ±$\pm 0.006$ 0.910 ±$\pm 0.002$
3 0.662 ±$\pm 0.027$ 0.562 ±$\pm 0.021$ 0.299 ±$\pm 0.023$ 0.705 ±$\pm 0.011$ 0.804 ±$\pm 0.006$ 0.915 ±$\pm 0.002$
$\frac{1}{\sqrt{d}}$ 0 0.567 ±$\pm 0.014$ 0.438 ±$\pm 0.033$ 0.181 ±$\pm 0.027$ 0.612 ±$\pm 0.008$ 0.646 ±$\pm 0.002$ 0.796 ±$\pm 0.002$
1 0.668 ±$\pm 0.017$ 0.597 ±$\pm 0.015$ 0.299 ±$\pm 0.019$ 0.692 ±$\pm 0.009$ 0.789 ±$\pm 0.003$ 0.904 ±$\pm 0.002$
2 0.676 ±$\pm 0.019$ 0.596 ±$\pm 0.014$ 0.312 ±$\pm 0.011$ 0.699 ±$\pm 0.009$ 0.805 ±$\pm 0.003$ 0.916 ±$\pm 0.002$
3 0.686 ±$\pm 0.025$ 0.602 ±$\pm 0.017$ 0.320 ±$\pm 0.013$ 0.708 ±$\pm 0.008$ 0.807 ±$\pm 0.005$ 0.919 ±$\pm 0.001$
Table 3 shows the results of our experiments, with the first row using the same hyperparameters as bhatt_end–end_2024 $(h_{\psi}=400,d_{\text{MLP}}=500)$. We also use these hyperparameters for UD and SciDTB, since our ablation study only concerns SemDP due to SynDP being less challenging. The lower half uses the best combinations for ADE (200, 100), CoNLL04 (400, 300), SciERC (300, 300), and ERFGC (400, 300), chosen based on mean performance across these four datasets.
Overall, normalizing biaffine scores provides an evident performance boost at all BiLSTM depths. In particular, for ERFGC we beat the state-of-the-art performance achieved by bhatt_end–end_2024 . Most of the performance gain is obtained by adding the first layer, with additional ones yielding diminishing returns. In other words, the performance boost provided by score normalization is highest in the absence of the implicit normalization provided by extra parameters ($L_{\psi}>0$). In general, the top performance obtained by using three BiLSTM layers and no score normalization can be matched or surpassed with a single BiLSTM layer, when using score normalization. Taking into account the lower values for $h_{\psi}$ and $d_{\text{MLP}}$, this represents a reduction in trained parameters of up to 85%.
@@
As laid out in Section 3, score variance tends to decay with deeper BiLSTM stacks. This in turn produces a converging trend as $L_{\psi}$ increases. When looking at Figure 3, this is especially evident for ERFGC, enEWT, and SciDTB, for which the beneficial effect of score normalization shrinks smoothly with higher values of $L_{\psi}$. Although the same cannot be said for ADE, CoNLL04, and SciERC, the performance boost is still statistically significant across all layer depths ($p<0.01$).
As regards SciERC, normalizing scores without any BiLSTM layers ($L_{\psi}=0$) produces a 23% increase in performance with a strong reduction in standard deviation. SciERC arguably has the hardest dependency graphs to parse, due to the little overlap between training and testing entities, which makes it difficult to leverage tag embeddings. In addition, it is characterized by complex semantic dependencies. Therefore, we reckon normalizing biaffine scores to be especially important when dealing with hard tasks. In particular, normalizing scores helps mitigate the higher variance of the model’s predictions for this challenging dataset. Conversely, the lack of score normalization exacerbates already uncertain predictions. For SciERC, this makes the trend observed in Figure 3 more unstable. Indeed, the performance first drops at $L_{\psi}=2$ and then increases once again at $L_{\psi}=3$, which does not happen for the other datasets.

### v2

@@ 4.1 Data
To train and evaluate our models, we use four SemDP datasets, [^1] We neglect ACE04 mitchell2005ace and ACE05 walker2006ace , two popular datasets for RE, due to their prohibitive cost. along with the 2.2 version of the Universal Dependencies English EWT treebank (enEWT) ud22 and SciDTB yang2018scidtb for SynDP. Table 1 summarizes datasets statistics and reports their entity and relation class annotations.
@@
As regards SemDP, ADE gurulingappa2012ade is a medical-domain dataset comprising reports of drug adverse-effect reactions. Each sample contains a single relation, “adverseEffect,” which links a disease to a drug. CoNLL04 roth2004conll04 contains news texts and is annotated with the classic named entities $e_{i}\in$ {per, org, loc} and relations $r_{j}\in$ {workFor, kill, orgBasedIn, liveIn, locIn}. ADE and CoNLL04 are characterized by relations which are not complex enough to form connected graphs of considerable size. SciERC luan2018scierc is a dataset compiled from sentences extracted from artificial intelligence literature. It is arguably the most challenging dataset used in this study, since most of the entities in the validation and testing partitions are not assigned an entity class. This makes it hard to train tag embeddings that can help to infer edges. ERFGC yamakata_english_2020 is a dataset comprising “flow graphs” parsed from culinary recipes. The semantic dependency graphs of these recipes are directed and acyclic, with a single root. Differently from bhatt_end–end_2024 , and as advised directly by yamakata_english_2020 (personal correspondence), we ignore the “-” edge labels present in the corpus, since they link discontinuous parts of phrasal verbs and are thus irrelevant. For ADE, CoNLL04, and SciERC we use the splits provided in bi_codekgc_2024 . [^2] https://drive.google.com/drive/folders/1vVKJIUzK4hIipfdEGmS0CCoFmUmZwOQV ERFGC is not available online; we obtained it by contacting the authors of yamakata_english_2020 .
@@
As regards SynDP, we use enEWT ud22 to compare directly against ji_graph-based_2019 ’s results. While many works evaluate SynDP on the Penn Treebank prasad-etal-2008-penn , it is closed-access and prohibitively expensive. Instead, we use SciDTB yang2018scidtb , a discourse analysis dataset comprising 798 abstracts extracted from the ACL Anthology. [^3] https://aclanthology.org It was processed for the syntax dependency parsing task using Stanza qi2020stanza . From these two datasets, we consider the xPOS tags (e.g., noun, preposition, determiner) and the syntactic dependencies (e.g., sentence root, object, adverbial modifier) to respectively be the entities and relations of the dependency graphs to infer.
In order to show that the normalization effect is language-independent, in Appendix A we also carry out SynDP experiments using Universal Dependencies nivre-etal-2020-universal datasets in six other languages. Finally, we also conduct experiments on three non-linguistic datasets: PCQM-Contact dwivedi2022lrgb , CIFAR10 Superpixel dwivedi2023benchmarking , and QM9 ramakrishnan2014qm9 . PCQM-Contact and QM9 are datasets which contain graphs of molecules, while CIFAR10 Superpixel contains superpixel clusters constructed from CIFAR10 images. We use these datasets to further show that the observed effects of normalization are not limited to linguistic dependency parsing, but are a general phenomenon of carrying out inference of fully connected graphs. We report the results of this additional experiment in Appendix B.
Table 1: Size of the train/dev/test splits and entity/relation classes for each dataset.
Data Entities Relations
ADE gurulingappa2012ade 2,563 / 854 / 300 disease, drug adverseEffect
CoNLL04 roth2004conll04 922 / 231 / 288 organization, person, location kill, locatedIn, workFor, orgBasedIn, liveIn
SciERC luan2018scierc 1,366 / 187 / 397 generic, material, method, metric, otherSciTerm, task usedFor, featureOf, hyponymOf, evaluateFor, partOf, compare, conjunction
ERFGC yamakata_english_2020 242 / 29 / 29 food, tool, duration, quantity, actionByChef, discontAction, actionByFood, actionByTool, foodState, toolState agent, target, indirectObject, toolComplement, foodComplement, foodEq, foodPartOf, foodSet, toolEq, toolPartOf, actionEq, timingHeadVerb, other
enEWT ud22 10,098 / 1,431 / 1,427 xPOS tags UD relations
SciDTB yang2018scidtb 2,567 / 814 / 817 xPOS tags UD relations
@@ 4.2 Model
$\textbf{h}_{i}^{tag}=$ $\phi(\textbf{x}_{i}),\quad\textbf{h}^{tag}_{i}\in\mathbb{R}^{d_{h}},$
$\textbf{y}_{i}^{tag}=$ $\text{Softmax}(\text{MLP}^{tag}(\textbf{h}_{i}^{tag})),\quad\textbf{y}_{i}^{tag}\in\mathbb{R}^{\mid T\mid},$
@@
In the parser, an optional $N$-layered BiLSTM $\psi$ produces new representations $\textbf{h}_{i}=\psi(\textbf{e}_{i}^{tag}\oplus\textbf{x}_{i})$, which are then projected into four different representations:
$\textbf{e}_{i}^{h}$ $=\text{MLP}^{(edge-head)}(\textbf{h}_{i}),$ $\textbf{e}_{i}^{d}=\text{MLP}^{(edge-dept)}(\textbf{h}_{i}),$
$\textbf{r}_{i}^{h}$ $=\text{MLP}^{(rel-dept)}(\textbf{h}_{i}),$ $\textbf{r}_{i}^{d}=\text{MLP}^{(rel-head)}(\textbf{h}_{i}).$
@@
$f(\textbf{x}_{1},\textbf{x}_{2};W)=\textbf{x}_{1}^{\top}W\textbf{x}_{2}+\textbf{x}_{1}^{\top}\textbf{b},$
$s_{i}^{edge}$ $=f^{(edge)}(\textbf{e}_{i}^{h},\textbf{e}_{i}^{d};W_{e}),$ $W_{e}\in\mathbb{R}^{d\times 1\times d},$
$s_{i}^{rel}$ $=f^{(rel)}(\textbf{r}_{i}^{h},\textbf{r}_{i}^{d};W_{r}),$ $W_{r}\in\mathbb{R}^{d\times|R|\times d},$
where $R$ is the set of relation classes, i.e., the possible labels applied to an edge. In Appendix C, following ji_graph-based_2019 , we extend the architecture and instead feed the representations produced by the BERT encoder into $L_{\gamma}\in\{0,1,2,3\}$ pairs of biaffine and graph attention network (GAT) veličković2018graph layers. This allows us to see the influence of normalization when using a $k$-hop parser.
Finally, in the decoder, the edge scores are used in conjunction with the relation representations $\textbf{r}_{i}^{h}$ and $\textbf{r}_{i}^{d}$ to obtain the final predictions. During training, we do greedy decoding, while during inference, we use Chu-Liu/Edmonds’ maximum spanning tree (MST) algorithm edmonds_optimum_1967 to ensure the predictions are well-formed trees. This is especially useful with big dependency graphs, since greedy decoding is more likely to produce invalid trees as size increases. When doing greedy decoding, an edge index (i.e., an adjacency matrix) $a_{i}=\arg\max_{j}s_{ij}^{edge}$ is produced by taking the argmax of the attention scores $s_{i}^{edge}$ across the last dimension. The edge index is then used to select which head relation representations $\mathbf{r}_{i}^{h}$ to use to calculate the relation scores $s_{i}^{rel}=f(\textbf{r}_{i}^{h},\textbf{r}_{i}^{d};W),\quad W\in\mathbb{R}^{d\times\mid R\mid\times d}$. The relations are then predicted as $r_{i}=\arg\max_{j}s_{ij}^{rel}$. When using MST decoding, edge and relation scores are combined into a single energy matrix where each entry represents the score of a specific head-dependent pair with its most likely relation type. This energy matrix is then used in the MST algorithm, producing trees with a single root and no cycles. Prior to energy calculation, edge scores and relation scores are scaled so that low values are squished and high values are increased, making the log softmax produce a hard adjacency matrix.
@@
$\mathcal{L}_{tag}$ $=-\frac{1}{|\mathcal{V}|}\sum_{i=1}^{|\mathcal{V}|}\sum_{t=1}^{|T|}y_{i,t}^{tag}\,\log p\bigl(y_{i,t}^{tag}\bigr),$
$\mathcal{L}_{edge}$ $=-\sum_{i,j=1}^{|\mathcal{V}|}\log p\bigl(y_{i,j}^{edge}=1\bigr),$
$\mathcal{L}_{rel}$ $=-\sum_{i=1}^{|\mathcal{V}|}\sum_{j=1}^{|\mathcal{V}|}\mathbbm{1}\bigl(y_{i,j}^{edge}=1\bigr)\sum_{\ell=1}^{|R|}y_{i,j,\ell}^{rel}\,\log p\bigl(y_{i,j,\ell}^{rel}\bigr),$
$\mathcal{L}$ $=\lambda_{1}\,\mathcal{L}_{tag}+\lambda_{2}\bigl(\mathcal{L}_{edge}+\mathcal{L}_{rel}\bigr).$
Losses are calculated based on the gold tags, edges, and relations. We set $\lambda_{1}=0.1$ and $\lambda_{2}=1$ as hyperparameters because the tagging task is much simpler than predicting the edges, since the same top performance is always achieved regardless of any other selected architecture hyperparameters. Following the usual approach for SynDP dozat_deep_2017 ; ji_graph-based_2019 ; jiang_entity-relation_2024 , when training on enEWT and SciDTB we use an oracle —the gold tags— and do not predict the POS tags ourselves. Since in this case we only focus on training the edge and relation classification tasks, we set $\lambda_{1}=0$.
@@ 4.3 Hyperparameters
We experiment with a range of hyperparameters for the encoder, tagger, and parser, as listed in Table 2. We use BERTbase devlin-etal-2019-bert as our pre-trained encoder, which we keep frozen throughout the whole training run in our main setting.
As regards the tagger, we set $L_{\phi}=1$ and $h_{\phi}=100$, as in bhatt_end–end_2024 , with weights initialized with a Xavier uniform distribution. In Appendix D.1, we ablate the $\phi$ BiLSTM and the use of the tag embeddings $\mathbf{e}^{tag}_{i}$ to assess their impact on overall performance. With relation to the parser, for all hyperparameter combinations of $\{N,h_{\psi},d_{\text{MLP}}\}$, we run our experiments by initializing its weights with a Xavier uniform distribution ($I_{par}\sim\mathcal{U}$) and no LayerNorm $\texttt{LN}_{\psi}$. In Sections˜D.2, D.3, D.4 and D.5, we conduct ablations over the parser hyperparameters indicated in Table 2.
Table 2: Main setting hyperparameter ranges. $\psi$ = Parser BiLSTM. $f$ = biaffine layer.
@@
MLP(edge) and $f^{(edge)}$ init. $I_{\text{par}}\sim\{\mathcal{U},\mathcal{N}\}$
@@
We train our models for $2k$ steps and evaluate on the development partition of each dataset every 100 steps. For enEWT and SciDTB, we train for $5k$ steps with $1k$-step validation intervals to make our results more comparable with the state of the art ji_graph-based_2019 . We apply early stopping at 30% of the total steps without improvement and choose the best model based on the top performance on the development split. In Appendix D.6, we extend the training to $10k$ steps and fully fine-tune a variety of small and large pre-trained language models to show time-wise test performance trends more clearly. We set the learning rate at $\eta=1\times 10^{-3}$ when the encoder is kept frozen. In all settings, including ablations, we use AdamW loshchilov2019adamw as the optimizer and a batch size of 8.
@@ 4.4 Evaluation
Following bhatt_end–end_2024 ; dozat_simpler_2018 ; jiang_entity-relation_2024 , we measure tagging and parsing performance on ADE, CoNLL04, SciERC, and ERFGC in terms of micro-averaged F1-measure. In addition, for enEWT and SciDTB we use unlabeled (UAS) and labeled (LAS) attachment score nivre2017universal . To corroborate the validity of our results, we train and evaluate each setting, including ablations, with five random seeds. We report mean and standard deviation for the F1, UAS, and LAS metrics, averaged over the five runs. To test the significance of our results, we use the one-tailed Wilcoxon signed-rank test woolson2005wilcoxon .
@@ 5 Results and discussion
Model $a$ NN ADE CoNLL04 SciERC ERFGC enEWT SciDTB
@@
Ours 11 0 0.541 $\pm 0.021$ 0.399 $\pm 0.024$ 0.147 $\pm 0.049$ 0.548 $\pm 0.010$ 0.559 $\pm 0.005$ 0.729 $\pm 0.004$
1 0.657 $\pm 0.011$ 0.556 $\pm 0.021$ 0.282 $\pm 0.009$ 0.676 $\pm 0.010$ 0.771 $\pm 0.006$ 0.892 $\pm 0.002$
2 0.667 $\pm 0.011$ 0.573 $\pm 0.025$ 0.273 $\pm 0.010$ 0.694 $\pm 0.010$ 0.796 $\pm 0.006$ 0.910 $\pm 0.002$
3 0.662 $\pm 0.027$ 0.562 $\pm 0.021$ 0.299 $\pm 0.023$ 0.705 $\pm 0.011$ 0.804 $\pm 0.006$ 0.915 $\pm 0.002$
$\frac{1}{\sqrt{d}}$ 0 0.567 $\pm 0.014$ 0.438 $\pm 0.033$ 0.181 $\pm 0.027$ 0.612 $\pm 0.008$ 0.646 $\pm 0.002$ 0.796 $\pm 0.002$
1 0.668 $\pm 0.017$ 0.597 $\pm 0.015$ 0.299 $\pm 0.019$ 0.692 $\pm 0.009$ 0.789 $\pm 0.003$ 0.904 $\pm 0.002$
2 0.676 $\pm 0.019$ 0.596 $\pm 0.014$ 0.312 $\pm 0.011$ 0.699 $\pm 0.009$ 0.805 $\pm 0.003$ 0.916 $\pm 0.002$
3 0.686 $\pm 0.025$ 0.602 $\pm 0.017$ 0.320 $\pm 0.013$ 0.708 $\pm 0.008$ 0.807 $\pm 0.005$ 0.919 $\pm 0.001$
Table 3 shows the results of our experiments for the labeled edge prediction task, with the first row using the same hyperparameters as bhatt_end–end_2024 $(h_{\psi}=400,d_{\text{MLP}}=500)$. We also use these hyperparameters for enEWT and SciDTB, since our ablation studies only concerns SemDP due to SynDP being less challenging. The lower half uses the best combinations for ADE (200, 100), CoNLL04 (400, 300), SciERC (300, 300), and ERFGC (400, 300), chosen based on mean performance across these four datasets. For brevity, the performance on the tagging and unlabeled edge prediction tasks are reported in Appendix E. Details on the used computational resources can be found in Appendix F.
Overall, normalizing biaffine scores provides an evident performance boost at all BiLSTM depths. In particular, for ERFGC we beat the state-of-the-art performance achieved by bhatt_end–end_2024 . Most of the performance gain is obtained by adding the first layer, with additional ones yielding diminishing returns. In other words, the performance boost provided by score normalization is highest in the absence of the implicit normalization provided by extra parameters ($N>0$). In general, the top performance obtained by using three BiLSTM layers and no score normalization can be matched or surpassed with a single BiLSTM layer, when using score normalization. Taking into account the lower values for $h_{\psi}$ and $d_{\text{MLP}}$, this represents a reduction in trained parameters of up to 85%.
Figure 3: Micro-F1 (SemDP) and LAS (SynDP) vs $N\in\{0,1,2,3\}$ at 2kk training steps. Red = norm; blue = raw. Performance increase with normalization is statistically significant ($p<0.01$).
@@
As laid out in Section 3, score variance tends to decay with deeper BiLSTM stacks. This in turn produces a converging trend as $N$ increases. When looking at Figure 3, this is especially evident for ERFGC, enEWT, and SciDTB, for which the beneficial effect of score normalization shrinks smoothly with higher values of $N$. Although the same cannot be said for ADE, CoNLL04, and SciERC, the performance boost is still statistically significant across all layer depths ($p<0.01$).
As regards SciERC, normalizing scores without any BiLSTM layers ($N=0$) produces a 23% increase in performance with a strong reduction in standard deviation. SciERC arguably has the hardest dependency graphs to parse, due to the little overlap between training and testing entities, which makes it difficult to leverage tag embeddings. In addition, it is characterized by complex semantic dependencies. Therefore, we reckon normalizing biaffine scores to be especially important when dealing with hard tasks. In particular, normalizing scores helps mitigate the higher variance of the model’s predictions for this challenging dataset. Conversely, the lack of score normalization exacerbates already uncertain predictions. For SciERC, this makes the trend observed in Figure 3 more unstable. Indeed, the performance first drops at $N=2$ and then increases once again at $N=3$, which does not happen for the other datasets.
Compared to existing approaches which do not scale the biaffine scores, fewer parameters can thus be trained to obtain similar performance. Therefore, our results indicate that parsers using raw biaffine scores are likely overparameterized, compared to the performance they could achieve by using score normalization.